# The Kalb-Ramond field and Gravitational Parity Violation

### v1

@@ VI Comments on constraints
The possibility of parity violation to be observed in gravitational wave detectors was suggested well over 10 years ago Alexander:2007kv ; Yunes:2010yf . The data currently being gathered from merger processes LIGOScientific:2014pky ; VIRGO:2014yos ; LIGOScientific:2020ibl ; KAGRA:2021vkt can in principle constrain parity violating parameters, although much more progress on this front will be made from third generation detectors due to their increased sensitivities Maggiore:2019uih ; Evans:2021gyd . Additionally, very recently, an analysis to constrain parity violation was carried out in Callister:2023tws using data from pulsar timing arrays NANOGrav:2023gor ; NANOGrav:2023hvm ; EPTA:2023fyk ; Xu:2023wog ; Reardon:2023gzh . The two primary effects of interest are velocity birefringence and amplitude birefringence. The former refers to right- and left-handed waves propagating with different frequency dependent phase shifts, the latter corresponding to an attenuation or amplification of one polarization over another. It is also the case that modified theories of gravity generally predict a propagation speed different from unity. In what follows, we will focus on propagation speed and amplitude birefringence.
@@ VI.1 Propagation speed
$-3\times 10^{-15}<c_{\text{gw}}-1<7\times 10^{-16}.$ (49)
The measurement was used to constrain parity violating effects in ghost-free scalar tensor models in Nishizawa:2018srh and in symmetric teleparallel gravity in Conroy:2019ibo . Using our result for the group velocity (47), this translates to a bound on the parameter $\tilde{\mathcal{V}}_{m}=\mathcal{V}_{m}/m_{p}$. If we take the dimensionless coupling to be $\xi\sim O(1)$, and use the upper bound on (49), then
$\mathcal{V}_{m}^{4}\Bigg{(}\frac{\mathcal{H}^{2}}{k^{2}m_{p}^{4}}\Bigg{)}<7\times 10^{-17}.$ (50)
This can be simplified in terms of the Hubble constant and proper wavenumber as
$\mathcal{V}_{m}\sqrt{\frac{H_{0}}{ka}}\lesssim 10^{-4}m_{p}.$ (51)
VI.2 Amplitude birefringence
[…]
$\mathcal{V}_{m}\lesssim 2.4\times 10^{8}\text{ GeV}.$ (52)
That the upper bound on the $\mathcal{V}_{m}$ parameter lies in the $O(100)$ PeV range is quite interesting. From a purely theoretical standpoint, it has been known for some time Wells:2004di that the PeV scale is an attractive range for supersymmetry (SUSY), admitting unified gauge couplings, disallowing flavor changing neutral currents, and predicting a Higgs mass between 125-155 GeV Bhattiprolu:2023lfh . It has also gained attention as a candidate energy scale for dark matter Berlin:2016vnh , partially motivated by the detection of near-PeV gamma rays by the Tibet $\text{AS}_{\gamma}$ collaboration TibetASgamma:2021tpz . Stringent constraints on PeV scale dark matter were derived in Maity:2021umk , and future data from Tibet $\text{AS}_{\gamma}$ is expected to further shed light on the viability of PeV scale dark matter (see also Cohen:2016uyg ; Bhattacharya:2019ucd ; Ishiwata:2019aet ; Blanco:2018esa for further relevant discussion of the gamma ray background).
@@ VII Discussion
Third generation gravitational wave detectors such as the Einstein Telescope Maggiore:2019uih and Cosmic Explorer Evans:2021gyd are anticipated to have sensitivities capable of probing beyond GR effects such as parity violation Califano:2023aji . Parity violation may also be observed in large scale structure Vanzan:2023cze ; Creque-Sarbinowski:2023wmb . Any observation definitively signalling parity violation would be a smoking gun for gravitational physics beyond General Relativity. It is therefore of great interest to explore various models which predict gravitational parity violation in anticipation of such data. In this work, inspired by recent attention on the Kalb-Ramond field as a dark matter candidate, we considered parity violation from a dimension-four operator coupling the dual Riemann tensor to the KR fields, $\mathcal{L}_{\text{int}}=-\tfrac{\xi}{2}\tilde{R}_{\mu\nu\rho\sigma}B^{\mu\nu}B^{\rho\sigma}$. For simplicity, we constructed a scenario where the KR fields can be treated as being constant. This will happen provided the KR field experiences a potential that satisfies $\frac{\partial V}{\partial B^{\mu\nu}}\big{|}_{B^{(0)}_{\mu\nu}}=0$ for $B^{(0)}_{\mu\nu}\neq 0.$ Two parameters $\mathcal{V}_{e},\mathcal{V}_{m}$, characterize the degrees of freedom of $B_{\mu\nu}.$ After writing down the field equations for the gravitational waves in the circular basis for arbitrary $\mathcal{V}_{m},\mathcal{V}_{e},$ we restricted to the case $\mathcal{V}_{e}=0$ before mapping the solution to the general parameterization presented in Jenks:2023pmk . Constraints from the GRB/gravitational wave signal GW170817 LIGOScientific:2017zic and the GWTC-3 KAGRA:2021vkt data set were considered, and the latter translated to the bound $\mathcal{V}_{m}\lesssim 10^{8}$ GeV. As we discussed in Sec. VI.2, this upper bound for dark matter being near the PeV scale is interesting from theoretical Wells:2004di ; Bhattiprolu:2023lfh and phenomenological TibetASgamma:2021tpz ; Maity:2021umk standpoints.
@@
We assumed a generic potential was responsible for the KR field settling into its background value, which triggers spontaneous Lorentz violation. As a result, (pseudo-) Nambu-Goldstone modes associated to the field fluctuations along the broken generators emerge Bluhm:2004ep ; Bluhm:2007bd ; Higashijima:2001sq ; Hernaski:2016dyk . Consequences of this spontaneous symmetry breaking were studied extensively in Altschul:2009ae , where the authors derived reasonable constraints on the Lorentz violating coefficients present in certain non-minimal curvature couplings (see, e.g. Cheng:2006us ; Carroll:2004ai ; Lambiase:2005kb ; Colladay:1998fq ; Kostelecky:1988zi ; Kostelecky:2008ts ; Kostelecky:2003fs ). However, such modes do not play an important role in our work here.
@@
It is of significant importance to consider the consequences of the interaction (6) in early universe physics. The KR field and more general $p$-forms in inflationary scenarios have been been studied in Starobinsky inflation Elizalde:2018now , $F(R)$ theories Elizalde:2018rmz , anisotropic inflation Almeida:2019xzt , and in general cosmological backgrounds BeltranAlmeida:2018nin . Each of these works assumed even parity and gauge invariance. Gauge invariance being preserved implies the KR field can be described in terms of one (pseudo)scalar degree of freedom, which is not an approach that we can take due to the non gauge invariance of (6). It is therefore interesting to explore the consequences of (6) in inflationary scenarios, in particular the deviations from GR for the primordial gravitational radiation, as well as consistency with other cosmological observables (for example, parity violation in the CMB was considered very recently in Gaztaaga2024:2401.08288v1 ). Allowing the KR fields to be fully dynamical opposed to frozen in a constant background value will be relevant in such scenarios. One approach would be to study the interactions with gravity in their polarization basis, as written in Capanelli:2023uwv . On the other hand, as mentioned in Section II, genuinely preserving cosmological isotropy when the KR field is in its background value should necessitate multiple fields, which may result in modifications to our results in this paper. We leave this and other explorations to future work.
@@ ACKNOWLEDGEMENTS
The authors thank Heliudson Bernardo, Leah Jenks, Evan McDonough, and Andrew Svesko for helpful comments and discussions. SA and TM are supported by the Simons Foundation, Award 896696.

### v2

@@ VI Comments on constraints
The possibility of parity violation to be observed in gravitational wave detectors was suggested well over 10 years ago Alexander:2007kv ; Yunes:2010yf . The data currently being gathered from merger processes LIGOScientific:2014pky ; VIRGO:2014yos ; LIGOScientific:2020ibl ; KAGRA:2021vkt can in principle constrain parity violating parameters, although much more progress on this front will be made from third generation detectors due to their increased sensitivities Maggiore:2019uih ; Evans:2021gyd . Additionally, very recently, an analysis to constrain parity violation was carried out in Callister:2023tws using data from pulsar timing arrays NANOGrav:2023gor ; NANOGrav:2023hvm ; EPTA:2023fyk ; Xu:2023wog ; Reardon:2023gzh . The two primary effects of interest are velocity birefringence and amplitude birefringence. The former refers to right- and left-handed waves propagating with different frequency dependent phase shifts, the latter corresponding to an attenuation or amplification of one polarization over another. It is also the case that modified theories of gravity generally predict a propagation speed different from unity. In what follows, we will focus on propagation speed.
@@ VI.1 Propagation speed
$3\times 10^{-15}<c_{\text{gw}}-1<7\times 10^{-16}.$ (41)
The measurement was used to constrain parity violating effects in ghost-free scalar tensor models in Nishizawa:2018srh and in symmetric teleparallel gravity in Conroy:2019ibo . Using our result for the group velocity (39), this translates to a bound on the parameter $\tilde{v}=v/m_{p}$. If we take the dimensionless coupling to be $\xi\sim O(1)$, and use the upper bound on (41), then
$\frac{v^{4}}{m_{p}^{4}}\Big(\frac{\mathcal{H}}{k}\Big)^{2}=\frac{v^{4}}{m_{p}^{4}}\Big(\frac{H_{0}}{ka}\Big)^{2}<5.6\times 10^{-15},$ (42)
where $H_{0}$ is the Hubble constant and $a=a(t)$ is the scale factor in cosmic time. Using the SH0ES result Riess:2021jrx for the Hubble constant $H_{0}\approx 73.04$ km/s/Mpc, we can write this as a function of redshift $z$ and the gravitational wave frequency $f$ as
$\frac{v}{m_{p}}\lesssim 4.4\times 10^{5}\Bigg(\frac{f/\text{Hz}}{1+z}\Bigg)^{1/2}.$ (43)
This illustrates that observations of lower frequency gravitational waves at high redshift would give the most meaningful constraint on $v$. This is particularly challenging, however, since the lowest frequency gravitational waves observed in pulsar timing arrays are from the stochastic GW background NANOGrav:2023gor . Since the source of the background is not yet known, redshift determination is not feasible. If the GW background is determined to be dominantly caused by the physics of inflation, (43) will be useful in constraining the energy scale of the KR fields [^7] We note that a very similar model including couplings to the Riemann tensor (opposed to its dual) and the Ricci tensor will indeed give an explicit constraint of $v\lesssim 10^{11}$ GeV ToAppear . .
@@ VII Discussion
Third generation gravitational wave detectors such as the Einstein Telescope Maggiore:2019uih and Cosmic Explorer Evans:2021gyd are anticipated to have sensitivities capable of probing beyond GR effects such as parity violation Califano:2023aji . Parity violation may also be observed in large scale structure Vanzan:2023cze ; Creque-Sarbinowski:2023wmb . Any observation definitively signaling parity violation would be a smoking gun for gravitational physics beyond General Relativity. It is therefore of great interest to explore various models which predict gravitational parity violation in anticipation of such data. In this work, inspired by recent attention on the Kalb-Ramond field as a dark matter candidate, we considered parity violation from a dimension-four operator coupling the dual Riemann tensor to the KR fields, $\mathcal{L}_{\text{int}}=-\tfrac{\xi}{2}\tilde{R}_{\mu\nu\rho\sigma}B^{\mu\nu}B^{\rho\sigma}$. For simplicity, we constructed a scenario where the KR fields can be treated as being constant. This will happen provided the KR field experiences a potential that satisfies $\frac{\partial V}{\partial B^{\mu\nu}}\big|_{B^{(0)}_{\mu\nu}}=0$ for $B^{(0)}_{\mu\nu}\neq 0.$ Two parameters $\mathcal{V}_{e},\mathcal{V}_{m}$, characterize the degrees of freedom of $B_{\mu\nu}.$ After writing down the field equations for the gravitational waves in the circular basis for arbitrary $\mathcal{V}_{m},\mathcal{V}_{e},$ we mapped the solution to the general parameterization presented in Jenks:2023pmk , which depends solely on $v=\sqrt{\mathcal{V}_{m}^{2}-\mathcal{V}_{e}^{2}}$ at lowest order in $\mathcal{V}_{m}/m_{p},\mathcal{V}_{e}/m_{p}$. Constraints from the GRB/gravitational wave signal GW170817 LIGOScientific:2017zic produced the bound given by (43).
@@
We assumed a generic potential was responsible for the KR field settling into its background value, which triggers spontaneous Lorentz violation. As a result, (pseudo-) Nambu-Goldstone modes associated to the field fluctuations along the broken generators emerge Bluhm:2004ep ; Bluhm:2007bd ; Higashijima:2001sq ; Hernaski:2016dyk . Consequences of this spontaneous symmetry breaking were studied extensively in Altschul:2009ae , where the authors derived reasonable constraints on the Lorentz violating coefficients present in certain non-minimal curvature couplings (see, e.g. Cheng:2006us ; Carroll:2004ai ; Lambiase:2005kb ; Colladay:1998fq ; Kostelecky:1988zi ; Kostelecky:2008ts ; Kostelecky:2003fs ; Amarilo:2023wpn ). However, such modes do not play an important role in our work here.
Since we only observe velocity birenfrigence, detecting the parity violation from this simple model relies heavily on multimessenger signals a lá (43), whereas amplitude birefringence can be detected without the need for such a signal. In either case, definitively detecting gravitational wave birefringence is a formidable task that naturally depends on the sensitivities of the detectors Moore:2014lga . Significant efforts have gone into studying detection prospects in the stochastic GW background Seto:2006hf ; Kato:2015bye ; Smith:2016jqs ; Callister:2017ocg ; Domcke:2019zls ; Omiya:2020fvw ; LIGOScientific:2020tif ; Wu:2021kmd ; Sato-Polito:2021efu ; Omiya:2023rhj ; Mentasti:2023gmg ; AnilKumar:2023hza ; Callister:2023tws , in addition to analysis focusing on late-time merger events. In the event that a conclusive observation of birefringence is made in, say, detections from merger processes, we will not explicitly know which scenario was responsible provided the signal’s properties can satisfactorily be described within known bounds on the model parameters. We expect that discerning the correct model will only be possible if its (beyond standard cosmology) predictions are verified at different scales and in different physical systems. For example, one would need precision measurements from next generation CMB observatories, and perhaps primordial gravitational wave precision properties, to discriminate between models that in some systems, predict the same physics.
@@
It is of significant importance to consider the consequences of the interaction (6) in early universe physics. The KR field and more general $p$-forms in inflationary scenarios have been been studied in Starobinsky inflation Elizalde:2018now , $F(R)$ theories Elizalde:2018rmz , anisotropic inflation Almeida:2019xzt , and in general cosmological backgrounds BeltranAlmeida:2018nin . Each of these works assumed even parity and gauge invariance. Gauge invariance being preserved implies the KR field can be described in terms of one (pseudo)scalar degree of freedom, which is not an approach that we can take due to the non gauge invariance of (6). It is therefore interesting to explore the consequences of (6) in inflationary scenarios, in particular the deviations from GR for the primordial gravitational wave spectrum as explored in Kahniashvili:2020jgm ; Li:2024fxy ; Ellis:2020uid , in addition to consistency with other cosmological observables (for example, parity violation in the CMB Gaztaaga2024:2401.08288v1 ). Allowing the KR fields to be fully dynamical opposed to frozen in a constant background value will be relevant in such scenarios. One approach would be to study the interactions with gravity in their polarization basis, as written in Capanelli:2023uwv . On the other hand, as mentioned in Section II, genuinely preserving cosmological isotropy when the KR field is in its background value should necessitate multiple fields, which may result in modifications to our results in this paper.
It will be interesting to see whether or not considering constraints from parity-even beyond GR models can tighten the upper bound on energy scale for the KR field, $v=\sqrt{\mathcal{V}_{m}^{2}-\mathcal{V}_{e}^{2}}$. This scenario also nontrivially affects the background evolution, which, in the context of early universe physics, can have a drastic affect on cosmological observables. A complete analysis will further shed light on the viability of non-minimal dimension-four couplings between the Riemann tensor and the KR field. We leave this and other explorations to future work.
@@ ACKNOWLEDGEMENTS
The authors thank Heliudson Bernardo, Tatsuya Daniel, Leah Jenks, Evan McDonough, and Andrew Svesko for helpful comments and discussions. SA and TM are supported by the Simons Foundation, Award 896696.